# Cryptanalysis of PLWE based on zero-trace quadratic roots

### v1

@@ 3.1. Attack developed by [36] and [37]
However, the sufficient condition for Algorithm 2 to work is that $|\Sigma|<q$ and since the look-up table must be only constructed once, for some parameter instances where building this table is still feasible, our attack will succeed, as proved in Proposition 3.7. For instance, for $f(x)=x^{316}+1672x^{314}+2x^{2}+103347$ and $q=100003$, the irreducible factor $x^{2}+1672$ divides $f(x)$ and $r:=\mathrm{ord}_{q}(-1672)=42$. For $\sigma=0.04$ we have
$|\Sigma|\leq\left(4\sqrt{\left\lfloor\frac{N^{*}}{r}\right\rfloor}\sigma+1\right)^{r}\approx 82469<q.$
Hence, $\frac{|\Sigma|}{q}<0.8247$ and Algorithm 2 predicts that the attack on $R_{q,0}\times R_{q}$ will succeed at least the $74\%$ of times with just $6$ samples and $90\%$ with $12$ samples.
@@ 4. Coding examples
    Obtaining a prime of the desired size such that $-1$ is not a quadratic residue $\mathbb{F}_{q}$.
@@
$N$ $62$ $102$
$\sigma$ $8$ $8$
ntests $20$ $20$
M $5$ $5$
Aver. exec. time (secs.) $\simeq 6.1$ $\simeq 16.75$
Aver. sampling time (secs.) $\simeq 0.19$ $\simeq 2.7$
The executions of the two examples gave no failures for the uniform distribution, but fails approximately $5\%$ to $10\%$ of the tests for the PLWE distribution. To understand why, keep in mind that Algorithm 2 is essentially probabilistic and requires that there exists a $g\in\mathbb{F}_{q}$ such that $(Tr(b(\alpha))-a(\alpha)g)/2\in\Sigma$ for all samples in the set of samples $C$; but this will occur only with probability $0.95$ by definition of $\Sigma$, as per Equation (3.3).
Regarding execution times, the examples have been executed and measured on our platform, a Virtual Box configured with 1 GB of main memory running over Intel CORE i5, at 3.4 GHz, which is rather slow. The execution time for creating samples from both distributions, namely uniform and PLWE, are very similar; therefore we just give one figure for the execution time, which is applicable to both of them.

### v3

@@ 3.1. Attack developed by [13] and [14]
However, the sufficient condition for Algorithm 2 to work is that $|\Sigma|<q$ and since the look-up table must be only constructed once, for some parameter instances where building this table is still feasible, our attack will succeed, as proved in Proposition 3.8.
Example 3.12.
Consider the polynomial
$f(x)=x^{256}+92882x^{254}+x^{128}+92882x^{126}+x^{2}+92882,$
and $q=100003$. The irreducible factor $x^{2}+92882$ divides $f(x)$, $r:=\mathrm{ord}_{q}(-92882)=6$. For $\sigma=0.04$ we have
$|\Sigma|\leq\left(4\frac{N^{*}}{r}\sigma+1\right)^{r}<7126<q.$
Hence, $\frac{|\Sigma|}{q}<0.07126$ and Proposition 3.8 predicts that the attack on $R_{q,0}\times R_{q}$ will succeed at least $81.62\%$ of times with just 55 samples and $99.993\%$ with $8$ samples. This example works with a root of significantly larger order as those considered in [13].
Example 3.13.
Consider the polynomial
$f(x)=x^{80}+x^{78}-x^{66}-x^{64}-2x^{48}-2x^{46}+2x^{34}+2x^{32}+x^{16}+x^{14}-x^{2}+262$
and $q=263$. The irreducible factor $x^{2}+1$ divides $f(x)$ and $r:=\mathrm{ord}_{q}(-1)=1$. For $\sigma=0.18$ we have
$|\Sigma|\leq\left(4\frac{N^{*}}{r}\sigma+1\right)^{r}<227<263<q.$
Hence, $\frac{|\Sigma|}{q}<0.863$ and Proposition 3.8 predicts that the attack on $R_{q,0}\times R_{q}$ will succeed at least $96\%$ of times with just $60$ samples. Notice that $q=263$ is close to byte size (251, 253). These byte-size primes are used in the LAC cryptosystem. We have chosen $N=80$ to simplify the computations, while LAC requires at least $N=128$. Moreover, in LAC, the modulus polynomial is cyclotomic. However, this example shows that our attack applies to instances where the parameters are relatively close to those used in practical scenarios.
Example 3.14.
Consider the polynomial
$f(x)=x^{32}+x^{30}+x^{2}+264,$
and $q=263$. The irreducible factor $x^{2}+1$ divides $f(x)$ and $r:=\mathrm{ord}_{q}(-1)=1$. For $\sigma=0.502$ we have
$|\Sigma|\leq\left(4\frac{N^{*}}{r}\sigma+1\right)^{r}<261<263<q.$
Hence, $\frac{|\Sigma|}{q}<0.9933$ and Proposition 3.8 predicts that the attack on $R_{q,0}\times R_{q}$ will succeed at least $87\%$ of times with $1000$ samples. Notice that in this example, the variance can be chosen to be considerably larger than in previous examples, also along the lines of practical scenarios.
@@ 4. Coding examples
    Obtaining a prime of the desired size such that 1-1 is not a quadratic residue $\mathbb{F}_{q}$.
@@
NN $62$ $102$
$\sigma$ 88 88
ntests 2020 2020
M 55 55
Average sampling exec. time $\simeq 4.97$ $\simeq 16.65$
Average test exec. time $\simeq 0.94$ $\simeq 3.02$
The executions of the two examples gave no failures for the uniform distribution, but fails approximately $5\%$ to $10\%$ of the tests for the PLWE distribution. To understand why, keep in mind that Algorithm 2 is essentially probabilistic and requires that there exists a $g\in\mathbb{F}_{q}$ such that $(\operatorname{Tr}(b(\alpha))-a(\alpha)g)/2\in\Sigma$ for all samples in the set of samples $C$.
Regarding execution times (given in seconds), the examples have been executed and measured on our platform, a Virtual Box configured with 1 GB of main memory running over Intel CORE i5, at 3.4 GHz, which is rather slow. The execution time for the test and for creating samples from both distributions, namely uniform and PLWE, are very similar; therefore we just give one figure for the execution time and one figure for the sampling time, which is applicable to both distributions. It is also apparent that sampling is by far the most time-consuming operation.